LBTI AO

02:43
  • Waiting for it to get dark enough.
2:59
  • Presetting
  • Bin 1, 692 Hz, Gains: 3.23
  • MOS ERROR!
  • Still bright out, haven't even hit 12 degree twilight
3:16
  • Here we go again. It went to bin2, probably because the sky background is darker.
  • Bin 2 1153 Hz Gain: 5.38 0.74 0.48
  • 20230530_032930
  • Seeing is "meh". 1.30"-1.50".
3:37
  • Did a successful "nudge".
4:29
  • Some weird offset issue between LUCI and the AO. Re-sent the preset.
  • Bin 2; 1117 Hz. Mag 11.7 Gains: 3.19 0.69 0.46
  • 20230530_043203
4:45
  • New Object
  • Bin 2; 898 Hz; 5.28 0.48 0.45 Mag 13.4
  • Need telemetry every 5 min.
  • 20230530_045320
  • 20230530_045801
    20230530_050403
    20230530_051001
  • 20230530_051533
  • 20230530_052001
  • 20230530_052528
  • 20230530_053103
  • 20230530_053530
  • 20230530_054002
  • 20230530_054503
  • 20230530_055000
  • 20230530_055502
  • 20230530_060004
6:03
  • ADSEC FAILURE, no apparent reason.
  • New preset, same object.
  • 20230530_061300
  • 20230530_061803
  • 20230530_062306
  • 20230530_062804
  • 20230530_063305
  • 20230530_063817
  • 20230530_064305
  • 20230530_064816
  • 20230530_065304
  • 20230530_065806
  • 20230530_070304
7:06
  • New preset.
  • Bin 3, 321 Hz; Gains: 6.03 1.80 Mag 16.2
  • 20230530_071352
8:03
  • ESM Preset
  • Bin 4, 200Hz, Gain: 3.67 1.23; mag 15.6
  • 20230530_080628

-- GregoryTaylor - 30 May 2023
Topic revision: r2 - 30 May 2023, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback